# Private From Whom? Minimal Information Leakage in Auctions

### v1

@@ 3.3 Pruning and Incentive Compatibility
We briefly comment on the incentive properties of protocols. If the choice function the designer seeks to implement is dominant-strategy incentive compatible, then it is without loss to restrict to protocols $P=(G,\bm{S})$ which are also dominant-strategy incentive compatible (that is, $\bm{S}$ is an equilibrium of $G$ in dominant strategies). Recall that $g(z)$ is the outcome at terminal history $z$ and $z(\bm{S})$ is the terminal history reached at strategies $S$.
@@
A choice function $\phi$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to truthfully reveal their type: for all bidders $i$, $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$,
@@
A protocol $P=(G,\{S_{i}(\theta_{i})\}_{i\in N})$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to follow the designer’s strategy recommendation: for all bidders $i$, $\theta_{i}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i},\sigma_{i}^{\prime}\in\Sigma_{i}$,
@@
In any pruned protocol, each history is reached on-path for some type, so deviations in the game $G$ are essentially deviations where a bidder pretends to be a different type. As such, incentive compatibility of the choice function is able to rule out all possible deviations.
@@
Protocols do not necessarily preserve Bayesian incentive compatibility, however. A protocol could provide certain bidders with pertinent information before they make a choice. As such, interim beliefs may be different under the protocol implementing the mechanism compared to its normal form game, so Bayesian incentive compatibility does not translate. This may provide another novel justification for dominant strategy incentive compatibility: Dominant strategy incentive compatibility is robust to the details of how a mechanism is actually run whereas Bayesian incentive compatibility is not. Furthermore, our impossibility theorem only gets stronger if Bayesian incentive compatibility is required as it restricts the set of potential protocols the designer may use. In our application to first-price auctions, for which the choice function $\phi^{FPA}$ is not dominant-strategy incentive compatible, we show that certain auctions are not privacy dominated by any protocol implementing $\phi^{FPA}$. This immediately implies that these auctions are not privacy-dominated by any incentive-compatible protocol implementing $\phi^{FPA}$.
4 Characterizations
4.1 Impossibility Theorem
We provide a condition on the choice function $\phi$ under which any minimally bidder-informative protocol must distinguish type profiles that are not pivotal to the outcome. We begin by stating a condition analogous to the Corners Lemma of chor1991boolean, which is sufficient for non-existence of a fully contextually private protocol.
Definition 11.
A choice function $\phi$ violates the corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if
$\phi(\theta_{i},\bm{\theta}_{-i})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})=\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})\neq\phi(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime}).$
Figure˜1 displays a violation of the corners condition, with $\alpha,\gamma\in X$ representing the choice function outcome $\phi(\bm{\theta})$ at type profile $\bm{\theta}$. To implement $\phi$, the auctioneer must distinguish the top-right corner, but need not distinguish between the other three type profiles.
$\alpha$$\alpha$$\alpha$$\gamma$$\theta_{i}$$\theta_{i}^{\prime}$$\bm{\theta}_{-i}$$\bm{\theta}_{-i}^{\prime}$
Figure 1: Corners Condition Violation
As we described in the model section, bidders are often unable to distinguish distinct outcomes (such as different outcomes when they lose the auction). As such, the equality between $\phi(\theta_{i},\bm{\theta_{-i}})$ and $\phi(\theta_{i},\bm{\theta_{-i}}^{\prime})$ can be weakened to equivalence under $\sim_{\Omega_{i}}$. This motivates our indistinguishable corners condition, corresponding to cases where bidder $i$ cannot distinguish between three of the outcomes at the corners, but the auctioneer must distinguish the final corner to implement the choice function.
Definition 12.
A choice function $\phi$ violates the indistinguishable corners condition for bidder $i$ at type profiles $\{\theta_{i},\theta_{i}^{\prime}\}\times\{\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\}$ if:
  1. 1.
    $\theta_{i},\theta_{i}^{\prime}$ should not be distinguished at $\bm{\theta}_{-i}$:
  2. 2.
    At $\theta_{i}$, $i$ should not distinguish between $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}$:
  3. 3.
    $\theta_{i},\theta_{i}^{\prime}$ needs to be distinguished at $\bm{\theta}_{-i}^{\prime}$:
[…]
Figure 2: Indistinguishable Corners Condition Violation
Note that a violation of the (standard) corners condition is also a violation of the indistinguishable corners condition as $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})=\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$ implies $\phi(\theta_{i},\theta_{j},\bm{\theta}_{-ij})\sim_{\Omega_{i}}\phi(\theta_{i},\theta_{j}^{\prime},\bm{\theta}_{-ij})$ since $\sim_{\Omega_{i}}$ is defined solely as a partition over the outcome space (and each outcome is trivially in the same partition as itself). In many cases, there are strictly more violations of the indistinguishable corners condition than the corners condition. In the first-price auction, there are no violations of the corners condition but many violations of the indistinguishable corners condition (Lemma 5). In the second-price auction, there are violations of the corners condition, but also distinct violations of the indistinguishable corners condition (Lemma 7).
Suppose the designer implements some choice function $\phi$ that violates the indistinguishable corners condition. The driving force behind our impossibility theorem is that to prevent bidders from learning information about other bidders’ types, the auctioneer must ask superfluous questions, thus learning unnecessary information. That is, any minimally bidder-informative protocol which implements $\phi$ must distinguish type profile $(\theta_{i},\bm{\theta}_{-i})$ from type profile $(\theta_{i}^{\prime},\bm{\theta}_{-i})$.
@@ 4.1 Impossibility Theorem
As $\phi(\theta_{i},\bm{\theta_{-i}})=\phi(\theta_{i}^{\prime},\bm{\theta}_{-i})$, the auctioneer learns information which is not necessary to compute the outcome. For intuition, consider Figure˜2. Suppose bidder $i$ has type $\theta_{i}$. If other bidders have type profile $\bm{\theta}_{-i}^{\prime}$, then the auctioneer must eventually ask bidder $i$ to distinguish type $\theta_{i}$ and $\theta_{i}^{\prime}$, in order for the auctioneer to determine the outcome. This means that if the auctioneer never asks bidder $i$ to distinguish $\theta_{i}$ and $\theta_{i}^{\prime}$, then bidder $i$ understands the type profile is not $(\theta_{i},\bm{\theta}_{-i}^{\prime})$. This is additional information that bidder $i$ obtains from the protocol being run. A auctioneer can provide each bidder strictly less information by running a sealed-bid protocol instead. In such a sealed-bid protocol, bidder $i$ cannot distinguish between $\phi(\theta_{i},\bm{\theta}_{-i})$ and $\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$ based on neither their observation about the outcome nor their experience in the protocol.
To show Lemma˜2, we use the following lemma formalizing the intuition above: bidder $i$ learns about $\bm{\theta}_{-i}$ if the auctioneer only asks them questions for certain values of $\bm{\theta}_{-i}$.
Lemma 3.
Consider $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}$. If $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$ and $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ then $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$ and $o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i})\neq o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$.
@@
(Lemma˜3) Because $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, there exists an information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ at which $S_{i}(\theta_{i})(I_{i})\neq S_{i}(\theta_{i}^{\prime})(I_{i})$. Because $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, at every information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, we have $S_{i}(\theta_{i})(I_{i})=S_{i}(\theta_{i}^{\prime})(I_{i})$. Hence there exists some information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By an identical argument, there exists some information set on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i})\neq o_{i}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. ∎
We may then use Lemma˜3 to prove Lemma˜2.
Proof.
@@
(Lemma˜2) By Lemma 1, in any minimally bidder-informative protocol, bidders learn nothing except what they observe of the outcome. Now suppose $G$ is minimally bidder-informative and implements $\phi$. Consider any $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ which violate the indistinguishable corners condition in Definition˜12. Because protocol $P$ implements $\phi$, we have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. By the indistinguishable corners condition, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$, and hence by $P$ being minimally bidder-informative, $o_{i}(\theta_{i},\bm{\theta}_{-i})=o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By the contrapositive of Lemma˜3, we conclude that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎
This now gives us a set of types who must be “unnecessarily” queried in any minimally bidder-informative protocol. This inherently conflicts with protecting bidders’ privacy from the auctioneer.
Theorem 1 (Impossibility).
  1. (i)
    $P$ produces no contextual privacy violations for bidder $i$.
  2. (ii)
    $P$ is minimally bidder-informative.
The intuition is straightforward, given Lemma˜2: any minimally bidder-informative protocol must learn unnecessary information about bidder $i$, meaning the privacy of bidder $i$ is compromised.
Proof.

### v2

@@ 8 Conclusion
This appendix comments on the incentive properties of protocols. If the choice rule the designer seeks to implement is dominant-strategy incentive compatible, then it is without loss to restrict to protocols $P=(G,\bm{S})$ which are ex-post incentive compatible (Nagel and Saitto, 2024). Recall that $g(z)$ is the outcome at terminal history $z$ and $z(\bm{S})$ is the terminal history reached at strategies $S$.
@@
A choice rule $\phi$ is dominant-strategy incentive compatible if it is a dominant strategy for each bidder to truthfully reveal their type: for all bidders $i$, $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i}$,
@@
A protocol $P=(G,\{S_{i}(\theta_{i})\}_{i\in N})$ is ex-post incentive compatible if it is a dominant strategy for each bidder to follow the designer’s strategy recommendation: for all bidders $i$, $\theta_{i}\in\Theta_{i},\bm{\theta}_{-i}\in\bm{\Theta}_{-i},\sigma_{i}^{\prime}\in\Sigma_{i}$,
@@
In any pruned protocol, each history is reached on-path for some type, so deviations in the game $G$ are essentially deviations where a bidder pretends to be a different type. As such, incentive compatibility of the choice rule is able to rule out all possible deviations.
@@
Protocols do not necessarily preserve Bayesian incentive compatibility, however. A protocol could provide certain bidders with pertinent information before they make a choice. As such, interim beliefs may be different under the protocol implementing the mechanism compared to its normal form game, so Bayesian incentive compatibility does not translate. This may provide another novel justification for dominant strategy incentive compatibility: Dominant strategy incentive compatibility is robust to the details of how a mechanism is actually run whereas Bayesian incentive compatibility is not. Furthermore, our impossibility theorem only gets stronger if Bayesian incentive compatibility is required as it restricts the set of potential protocols the designer may use. In our application to first-price auctions, for which the choice rule $\phi^{FPA}$ is not dominant-strategy incentive compatible, we show that certain auctions are not privacy dominated by any protocol implementing $\phi^{FPA}$. This immediately implies that these auctions are not privacy dominated by any incentive-compatible protocol implementing $\phi^{FPA}$. For additional discussion, see Fadel and Segal (2009) investigating the communication costs of requiring incentive compatibility in extensive-form protocols.
Appendix B: Omitted Proofs
Impossibility Theorem
PROOF OF LEMMA 1
Proof.
(Lemma˜1) Because $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$, there exists an information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ at which $S_{i}(\theta_{i})(I_{i})\neq S_{i}(\theta_{i}^{\prime})(I_{i})$. Because $(\theta_{i},\bm{\theta}_{-i})\sim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$, at every information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, we have $S_{i}(\theta_{i})(I_{i})=S_{i}(\theta_{i}^{\prime})(I_{i})$. Hence there exists some information set $I_{i}$ on the path of play for $(\theta_{i},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i},\bm{\theta}_{-i})$, and therefore $o_{i}(\theta_{i},\bm{\theta}_{-i})\neq o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By definition, $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By an identical argument, there exists some information set on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$ which is not on the path of play for $(\theta_{i}^{\prime},\bm{\theta}_{-i})$, and therefore $(\theta_{i}^{\prime},\bm{\theta}_{-i})\nsim_{(P,i)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. The second sentence in the lemma follows from the contrapositive. ∎
PROOF OF THEOREM 1
Proof.
The proof proceeds in three steps. First, we characterize when bidder information leakages occur. Second, we use this to show that contextual privacy violations must occur if bidder information leakages do not. Finally, we connect these results to the statement of Theorem 1.
@@
(Lemma˜7) In any minimally bidder-informative protocol, bidders learn nothing except what they observe of the outcome. Now suppose $G$ is minimally bidder-informative and implements $\phi$. Consider any $\theta_{i},\theta_{i}^{\prime}\in\Theta_{i}$ and $\bm{\theta}_{-i},\bm{\theta}_{-i}^{\prime}\in\bm{\Theta}_{-i}$ which violate the indistinguishable corners condition in Definition˜3. Because protocol $P$ implements $\phi$, we have $(\theta_{i},\bm{\theta}_{-i}^{\prime})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i}^{\prime})$. By the indistinguishable corners condition, $\phi(\theta_{i},\bm{\theta}_{-i})\sim_{\Omega_{i}}\phi(\theta_{i},\bm{\theta}_{-i}^{\prime})$, and hence by $P$ being minimally bidder-informative, $o_{i}(\theta_{i},\bm{\theta}_{-i})=o_{i}(\theta_{i},\bm{\theta}_{-i}^{\prime})$. By the contrapositive of Lemma˜1, we conclude that $(\theta_{i},\bm{\theta}_{-i})\nsim_{(P,A)}(\theta_{i}^{\prime},\bm{\theta}_{-i})$. ∎